Grade upto college level Physical Chemistry

75.2 g of C6H5OH (phenol) is dissolved in a solvent of Kf = 14. If the depression in freezing point is 7 K then find the % of phenol that dimerises.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
2C6H5OH ⇌ (C6H5OH)2
Initial no. of moles 1 0
No. of moles at equilibrium 1 - ∝ ∝/2
Total number of moles at equilibrium = 1 - ∝ + ∝/2 = 1 - ∝/2
Tf = iKf* (molality)
⇒ 7 = 14 * 75.2/94 * (1 - ∝/2) [weight of phenol = 75.2g mol. wt of phenol = 94]
∴ ∝ = 0.75
So the percentage of phenol that dimerises = 75%.
ALTERNATIVESOLUTION :
Phenol dimerises in the solvent (organic) as :
2C6H5OH ⇌ (C6H5OH)2
1 mol 0 mol (Initial)
1 - ∝ mol ∝/2 mol (Equilibrium)
Van’t Hoff factor, i = 1 - ∝ + ∝/2 = 1 - ∝/2
Where ∝ is the degree of dimerization.
Tf(obs) = Kf * molality * i
Or 7 = 14 * m * (1 - ∝/2)
NOTE : In the given problem, amount of the solvent used to dissolved 75.2 g of phenol is not mentioned. Hence molality m, can not be calculated and so also the value of ∝.
If amount of solution is presumed to be 1 kg (= 100g) then,
M = 75.2/94 = 0.8 (mol. mass of phenol = 9.4)
And 7 = 14 * 75.2/94 * (1 - ∝/2)
Or ∝ = 0.75 = 75%

Thanks
Deepak patra
askIITians Faculty